Daily sales of lemonade and ice cream at your new stand depend on the days temperature as follows: Lemonade Ice Cream Probability Sales $80 $100 Sales $20 $100 $180 Warm Hot Scorching .40 40 20 $120 a. Calculate the expected daily lemonade sales. b. Calculate the expected daily ice cream sales. expected dans ween daily ice c. Calculate the covariance between daily ice cream sales and daily lemonade sales.
Answer #1

(a) Expected Daily Lemonade Sales = Probabilityofl × Salesofl = 0.4 x 80 + 0.4 x 100 + 0.2 x 120 = $ $ 96

(b) Expected Daily Ice-Cream Sales = 0.4 x 20 + 0.4 x 100 + 0.2 x 180 = $ 84

(c) Covariance = [ 0.4 x (80 - 96) x (20 - 84) + 0.4 x (100 - 96) x (100 - 84) + 0.2 x (120 - 96) x (180 - 84) ] ^(1/2) = [896]^(1/2) = 29.93 %